# Evolutionary Hysteresis: Cycling about in a Rugged Landscape

### v1

@@ II Results
To introduce epistasis into a haploid two-locus two-allele Wright-Fisher model (e.g. $a_{1}b_{1},a_{1}b_{2},a_{2}b_{1},a_{2}b_{2}$), we define a simple fitness landscape
@@
where we assign a relative fitness benefit $s$ to haplotypes containing alleles of type-1, with an epistatic correction $J$. Such constructions are standard in population genetics and allow for precise control of evolutionary parameters [karlin_general_1975]. The case $s<J$ is known as reciprocal sign epistasis, where the “mixed” haplotypes ($a_{1},b_{2}$) or $(a_{2},b_{1})$ is less fit than the “pure” haplotypes ($(a_{1},b_{1})$ and $(a_{2},b_{2})$). Reciprocal sign epistasis may be caused by functional and physical incompatibilities of biomolecules, such as charge mismatches of amino acids that disrupt protein folding or enzymatic sites [starr_epistasis_2016, domingo_causes_2019, kvitek_reciprocal_2011].
A standard diffusion approximation (Supplementary Information, SI.1) yields a Fokker-Planck equation governing the probability density $P(p_{a},p_{b})$ of frequencies $p_{a},p_{b}$ of haplotypes $a_{1},b_{1}$ (that of others are $p_{a}^{\prime}=1-p_{a}$ and $p_{b}^{\prime}=1-p_{b}$),
@@
where $A(p_{a},p_{b})$ and $B(p_{a},p_{b})$, are the advection and diffusion coefficients, and depend explicitly on the model parameters $s$, $J$, mutation rate $u$, and population size $N$. From Eq.1 and Eq.2, one obtains a stationary state of the population distribution for a parameter set $(s,J,u,N)$ of the form $P_{st}(\vec{p})\propto e^{-2N\psi(\vec{p})}$, where the $\psi(\vec{p})$ is called the fitness potential and is given by
@@
Figure 1: Hysteresis in the two-locus two-allele Wright-Fisher model with epistasis. (a) The effective fitness potential of an epistatic population (Eq. 4) turns from a single to a double-well form when there is significant reciprocal sign epistasis $(s<J)$ for mixed haplotypes. (b) In our simulations, allele frequencies exhibit a hysteresis loop as we sweep selection strength back and forth ($J=0.2$, red circles). The hysteresis loop area vanishes in the absence of epistasis ($J=0$, black circles). In (b), (c) The form of the curves and hysteresis area matches our analytical formulas (Eq. S2) for populations as small as $N=500$ (solid curves). (d) The demographic noise in smaller populations allow random jumps from one hysteresis branch to the other (i.e. between two potential minima). We determine this escape rate $\kappa_{\textrm{esc}}$ from the lower to upper stable allele state via simulations (circles) and analytical theory (Eq. 5, solid curves). (e) The phase diagram measuring the distance between the potential wells (blue to yellow). The white region is where the potential becomes single welled and hysteresis vanishes.
Due to the symmetric nature of Eq. 1 the stable points of $\psi$ fall along the line $p=p_{a}=p_{b}$, and the system can be reduced to one dimension, Eq.4.
@@
This fitness potential shows us how varying $s$ can cause a transition between a mono-stable and bistable system (Fig.1a,b). Because the height of the epistatic fitness barrier is set by $J$, reversing the sign of $s$ does not immediately erase the previous state, thus an evolving system “remembers” its local minimum and retains this state even if it no longer optimally fit. The red curve in Fig.1b is the characteristic hysteresis loop that is common to many history dependent systems. Fig.1e shows the degree of bistability (as quantified by the distance between wells) as a function of epistasis strength and mutation rate. Hysteresis does not occur in the white zone.
To examine the accuracy of the effective fitness potential to predict evolutionary outcomes of epistatic populations, we simulated the evolution of a population of $N$ governed by a Wright-Fisher process with mutation and epistatic selection defined by Eq.1. The cycling of the selection pressure $s$ resembles many common scenarios in nature, such as seasonal oscillations of light, temperature, nutrient and water availability [yu_distinct_2022].
From these results, we can evaluate the quality of the diffusion approximation as a function of $N$ and $u$. The diffusion approximation accurately describes the roughly linear relationship between epistatic strength and hysteresis area (Fig.1c, Eq.S2), and holds for populations as small as $N=500$.
@@
We next obtain the rate at which a population escapes over an epistatic fitness barrier. This rate measures the population’s adaptive mobility or evolvability: the likelihood of breaking free from an epistatic constraint to discover a new, fitter region of the landscape. Stronger epistasis or larger populations suppress this rate, rendering adaptation sluggish even when environmental conditions favor transition [weissman_rate_2009]. We use Eyring-Kramer theory to obtain the barrier escape rate $\kappa_{esc}\propto\sqrt{U(x)^{\prime\prime}(x_{0})|U^{\prime\prime}(x_{s})|}\exp[U(x_{0})-U(x_{s})]$, where $x_{0}$ and $x_{s}$ are the location of the minima and barrier of a generalized potential function $U(x)$ [eyring_activated_1935, kramers_brownian_1940]. Recent work has derived the correct expression for the Kramers’ rate for cases with state-dependent noise [rosas_kramers_2016], which we use to calculate the escape rate over an epistatic barrier. The scaling of the escape rate with $s,J,N,$ and $u$ is given by Eq.5, with a negligible prefactor (SI.2).
@@
Fig.1d shows how this rate decreases exponentially with increasing population size. For large populations, barrier crossing events become exceedingly rare, in agreement with previous results on stochastic tunneling [weissman_rate_2009, altland_rare_2011], but now with explicit dependence on the selection and epistasis parameters $s$ and $J$. As the selection strength increases, escape events become much more common.
These results suggest that environmental conditions and demographic constraints jointly determine whether populations remain trapped behind epistatic barriers or traverse them.
@@
To do so, we simulated the evolutionary trajectories of our epistatic two-locus two-allele population subject to a selective force with a cyclic component with amplitude $A$ and frequency $\Omega$ (to yield hysteresis) and a random component $\eta(t)$ (that will modulate the hysteresis),
@@
We observe that noise anneals hysteresis by enabling random jumps over epistatic barriers: Epistatic hysteresis curves (2a, blue), in the presence of moderate noise (red), approach non-epistatic ones (black), indicating enhanced evolvability (Fig.2a). However, as the noise is increased further, the population is unable to keep up with an environment fluctuating too wildly, and will suffer a fitness shortfall. Fig.2b shows the sweet spot (blue markers) between evolvability and unpredictability. The fitness shortfall $\bar{\varepsilon}$ is defined as the deviation from the ideal fitness $\bar{\omega}^{*}(s(t))$ (the fitness, had there been no epistatic obstruction, i.e. with $J=0$ substituted in Eq.3).
@@
where $\bar{\omega}(t)$ is the instantaneous mean fitness of the population at time $t$ (Eq. 3), and $T$ is the averaging time over many cycles (cf. Methods).
As expected, the optimal value of noise is a function of the ruggedness of the fitness landscape, which in our model is quantified by $J$ (Fig.S2). The higher the epistatic barrier, the larger the noise the population will benefit from.
@@
Figure 3: Exploring dynamical hysteresis. A single-well potential well moving back and forth quick enough will have the effect of a double well, leading to similar hysteresis effects, remarkably, even in the absence of epistasis ($J=0$). (a) As we increase oscillation frequency from $\Omega/u=0.01$ (black) to $\Omega/u=1.0$ (red), we observe the emergence of dynamic hysteresis, and the bistability that underlies it. (b) We use the probability density of alleles (inset of (a)) to fit an effective fitness potential via Eq. 4. As we increase the cycling frequency, an effective double-well structure emerges, mimicking the role of epistasis. (c) Regardless of the true value of epistasis (black: $J_{\text{sim}}=0$, blue: $J_{\text{sim}}=0.1$, red: $J_{\text{sim}}=0.2$), we can fit an “inferred effective epistasis” $J_{\text{fit}}$ to the emergent effective potential. For $J=0$ we observe a phase transition where hysteresis emerges once the cycling rate exceeds the evolutionary relaxation rate $(\omega/u)_{\textrm{crit}}$. This transition softens and gradually vanishes with larger values of true epistasis $J_{\text{sim}}$. (d) A phase diagram for $J=0$ shows the interplay between selection noise $\sigma$ and cycling rate, and delineates the regime where dynamical hysteresis emerges.
@@
The two-locus Wright-Fisher model revealed how hysteresis emerges as a consequence of epistasis. However, it is not obvious whether hysteresis is a sensitive phenomena specific to this simple model, or it is a robust, generic property exhibited realistic genomes where thousands of genes couple together through a complex network of epistatic interactions. For example, should we expect a large number of random (positive and negative) interactions $J_{ij}$ between gene pairs $i,j$ to “cancel out”, and wash away hysteresis? Moreover, will the physical, structural properties of a protein, which might not be simply a function of pairwise interactions, exhibit similar hysteresis loops?
We answer both questions by demonstrating the existence of hysteresis loops, first, in a complex disordered L-locus K-allele model, and then, by running biophysically realistic molecular evolution simulations, where residue–residue interactions arise from the biophysics of protein folding and structural flexibility (Fig.4a,b).
For the L-locus K-allele model, the epistatic interaction matrix $J_{ij}^{\alpha,\beta}$ is populated from a Gaussian with mean $\langle J\rangle$ and variance $\delta J$, corresponding to increased ruggedness of the fitness landscape. The fitness potential for this more complex model is obtained through the same procedure outlined above (SI.3),
@@
Having established that hysteresis emerges generically in both minimal and disordered epistatic systems, and out in the wild, we next ask whether the shape and size of the loop can serve as a probe to measure the evolutionary parameters of a real system. We show that by performing cyclic selection empirically, one can gain insights into $K$, $L$, $\langle J\rangle$, $\delta J$ in real systems (Fig.6b, Fig.S4).
To map the loop morphology to underlying evolutionary parameters, we expand the effective potential $\psi$ (Eq. 8) around the symmetric state $p_{li}=1/K$ in terms of the mean allelic deviation $m=\frac{1}{L}\sum_{l,i}(p_{li}-1/K)$, obtaining a coarse-grained potential $\phi(m)=a_{2}m^{2}+a_{4}m^{4}-hm$, with coefficients $a_{2}\simeq\frac{u}{K-1}\chi_{K}-\frac{1-u/K}{K-1}\lambda_{2}(J,L)+\alpha\,\delta J^{2}$ and $h\simeq\frac{1-u/K}{K-1}\lambda_{1}(L)s$, where $\chi_{K}$ arises from the entropic curvature of the mutation term and $\lambda_{1,2}$ describe the additive and pairwise contributions to $\ln\bar{\omega}(p)$. Hysteresis occurs when $a_{2}<0$, i.e., when the epistatic stabilization $\lambda_{2}(J,L)$ dominates mutation and disorder, and the barrier height between the wells approximately scales as $\Delta\phi\sim a_{2}^{2}/(4a_{4})$.
@@
In principle, the full shape of the loop, including branch curvature, asymmetry, and minor-loop closure, should contain enough information to reconstruct not only coarse descriptors such as barrier height and disorder, but also finer details of the fitness landscape and the epistatic network. Solving this formidable inverse problem would require incorporating the complete loop trajectory rather than summary quantities such as area or squareness, and developing analytic or machine-learning surrogates that map loop morphologies to landscape statistics. This approach could turn hysteresis measurements into a new quantitative probe of epistasis networks and fitness landscapes, but its realization lies beyond the scope of the present work.
@@ IV Methods
Here we describe the protocol used to simulate hysteresis loops for the epistatic Wright-Fisher model. Starting with a population with equal proportions of the four haplotypes, we evolved the population for $1000$ generations with fixed values of $s$ and $J$. We varied $s$ in a quasi-static cycle every $1000$ generations from $0.5$ to 0.5-0.5 and back to $0.5$ to complete one evolutionary cycle. For Fig.1c, this scheme was performed independently for $N=(500,10000)$, $u=0.001$, and 20 values of $J$ linearly spaced from 0 to $0.2$. We simulated $48$ independent replicate trajectories for each parameter set $(N,J,u)$, resulting in $18000$ total trajectories for analysis. For each trajectory, we computed the area between the upper ($s\rightarrow 0^{+}$) and lower ($s\rightarrow 0^{-}$) stationary points over a full selection cycle, which we call the hysteresis area. The stationary points were defined as the average frequency of the allele under selection in the last $250$ generations of a given leg of fixed selection strength $s$.
@@
The simulation results presented in Fig.1d were obtain as follows. We performed simulations of a one-dimensional Ornstein-Uhlenbeck process with advection and diffusion terms given by the terms $A(p)$ and $B(p)$, respectively (Eq. S1), which are functions of the model parameters $s,J,u$ and $N$. Starting from the local fitness maximum $p_{0}$, the stochastic variable $p(t)$ was allowed to evolve until it reached the barrier crossing point given by $p_{s}$, and the time of first crossing was recorded. We performed 1000 independent trials of this process for a range of population sizes N=500N=500-$10000$ and selection strengths $s=0.2,0.225,0.25$, with $J=0.5$. We report the average rate of escape $\kappa_{\textrm{esc}}=1/\langle T_{\textrm{esc}}\rangle$ with 95% confidence intervals determined by propagation of errors.
@@
The data presented in Fig.2a were obtained with simulation parameters $A=0.5,\Omega=0.001$, $\sigma=0,0.1371$ and $J=0,0.4$, and depict the average allele frequency of 48 independent replicate simulations. For the results presented in Fig.2b, we simulated 48 independent replicate trajectories of the epistatic Wright-Fisher model starting from equal haplotype populations for model parameters J=0J=0-$0.4$, $\sigma=0$-$0.1$, u=0.0001u=0.0001-$0.01$, $N=500$, $A=0.5$, and $\Omega=0.001$. This resulted in $19200$ total trajectories for analysis. The quantity $T$ in Eq. 7 represents the total number of selection stages $s(t)$, which was chosen to complete one full hysteresis cycle for a given cycling rate $\Omega$. The results presented in Fig.3 were obtained by 48 independent replicate trajectories for the following parameters: $J=0,0.03,0.1$, $N=500$, $u=0.01$, selection amplitude $A=0.5$, cycling rate $\Omega=0.0002$-$0.02$, and selection noise $\sigma=0$-$0.1$. The resulting allele frequencies around $s=0$ were fit to the effective fitness potential model by non-linear least squares optimization using the SciPy package with Python 3.9. Uncertainty in the fit parameters were estimated as the 95% confidence interval of 500 non-parametric bootstrap simulations.
@@
We performed simulations of the $L$-locus $K$-allele Wright-Fisher model (Eq. 8) with disordered epistatis to investigate the link between epistatic interaction networks and macroscopic hysteresis in long evolutionary cycles. The simulations were performed in a similar manner as for the two-locus two-allele investigations, involving slow cycling of selection pressure with fixed epistatic interaction parameters. First, we assigned single-site fitness parameters $s_{i}^{\alpha}$ to each allele type $i$ at site $\alpha$, with $s_{i}^{\alpha}$ ranging from 0 to maximum (or minimum) fitness value $s\in(-2,2)$. The epistatic parameters $J_{ij}^{\alpha,\beta}$ were generated randomly from a Gaussian distribution $\mathcal{N}(\langle J\rangle,\delta J)$ with mean $\langle J\rangle$ and variance $\delta J$. The simulations were run in sequential stages of fixed $s$ for 1000 generations each before moving on to the next selection target. At the end of each evolutionary stage, we calculated the response of the population by Eq. 9.
@@
To simulate the evolution of protein sequences under selective pressure and epistasis, we modified the Protein Evolution Parameter Calculation (PEPC, [piszkin_extremophile_2021]) to include protein sequence flexibility prediction using the MEDUSA prediction software [vander_meersche_medusa_2021]. This method classifies amino acid residues by contribution to local structural flexibility/rigidity. Since the MEDUSA tool was trained using homologous sequences and physico-chemical properties of amino acids, it necessarily incorporates epistatic interactions between residues. Fig.4a depicts the procedure for evolving protein sequences in the modified PEPC method. Every generation, a candidate amino acid substitution is determined from a BLAST query against the Pfam database and the BLOSUM80 matrix [piszkin_extremophile_2021]. Then, a MEDUSA prediction is run on the candidate mutated sequence to assign a confidence level for the placement of each residue into one of five flexibility classes. To increase throughput, the HHblits MSA construction step of the MEDUSA prediction was performed only if the sequence had diverged by more than 80% sequence identity from its sequence at the last MSA construction, otherwise the most recent MSA was used. The predicted global flexibility $F$ of the protein sequence was calculated from the MEDUSA results by Eq. 10, where $c_{l,a}$ is the confidence that residue $l$ belongs to the MEDUSA flexibility class $a$, and $n$ is the length of the protein.
@@
Here, the classes range from least flexibility class 1 to most flexible class 5. The global flexibility of the candidate sequence is compared to the parent sequence and the mutation is accepted with a probability determined by a Metropolis-Hasting criterion defined by Eq. 11.

### v2

@@ Abstract
Keywords: fitness landscape, epistasis, hysteresis, protein evolution, microbial evolution
Classification: Biological Sciences, Evolution
Significance Statement: Evolving populations are faced with adapting to changing environments, while interactions between genetic components constrain available evolutionary pathways. This leads to hysteresis, a history-dependence of evolution. We theoretically, computationally, and empirically characterize evolutionary hysteresis in simple and complex population genetic models, protein evolution with disordered epistasis, and real-world seasonal microbial evolution. Our findings demonstrate how hysteresis may serve to probe complex features of rugged fitness landscapes.
@@ II Results
To introduce epistasis into a two-locus two-allele Wright-Fisher model (e.g. $a_{1}b_{1},a_{1}b_{2},a_{2}b_{1},a_{2}b_{2}$), we define a simple fitness landscape
@@
where we assign a relative fitness benefit $s$ to genotypes containing alleles of type-1, with an epistatic correction $J$. Such constructions are standard in population genetics and allow for precise control of evolutionary parameters [karlin_general_1975]. The case $s<J$ is known as reciprocal sign epistasis, where the “mixed” genotypes ($a_{1},b_{2}$) or $(a_{2},b_{1})$ is less fit than the “pure” genotypes ($(a_{1},b_{1})$ and $(a_{2},b_{2})$). Reciprocal sign epistasis may be caused by functional and physical incompatibilities of biomolecules, such as charge mismatches of amino acids that disrupt protein folding or enzymatic sites [starr_epistasis_2016, domingo_causes_2019, kvitek_reciprocal_2011].
A standard diffusion approximation (Supporting Information, SI. I) yields a Fokker-Planck equation governing the probability density $P(p_{a},p_{b})$ of frequencies $p_{a},p_{b}$ of genotypes $a_{1},b_{1}$ (that of others are $p_{a}^{\prime}=1-p_{a}$ and $p_{b}^{\prime}=1-p_{b}$),
@@
where $A(p_{a},p_{b})$ and $B(p_{a},p_{b})$, are the advection and diffusion coefficients, and depend explicitly on the model parameters $s$, $J$, mutation rate $u$, and population size $N$. From Eqn.1 and Eqn.2, one obtains a potential solution of the stationary state of the population distribution for a parameter set $(s,J,u,N)$ of the form $P_{st}(\vec{p})\propto e^{-2N\psi(\vec{p})}$, where the $\psi(\vec{p})$ is called the fitness potential and is given by
@@
Figure 1: Hysteresis in the two-locus two-allele Wright-Fisher model with epistasis. (a) The effective fitness potential of an epistatic population (Eqn.4) turns from a single to a double-well form when there is significant reciprocal sign epistasis $(s<J)$ for mixed genotypes. (b) In our simulations, allele frequencies exhibit a hysteresis loop as we sweep selection strength back and forth ($J=0.2$, red circles). The hysteresis loop area vanishes in the absence of epistasis ($J=0$, black circles). In (b), (c) The form of the curves and hysteresis area matches our analytical formulas (Eqn.S2) for populations as small as $N=500$ (solid curves). (d) The demographic noise in smaller populations allow random jumps from one hysteresis branch to the other (i.e. between two potential minima). We determine this escape rate $\kappa_{\textrm{esc}}$ from the lower to upper stable allele state via simulations (circles) and analytical theory (Eqn.5, solid curves). (e) The phase diagram measuring the distance between the potential wells (blue to yellow). The white region is where the potential becomes single welled and hysteresis vanishes.
Due to the symmetric nature of Eqn.1 the stable points of $\psi$ fall along the line $p=p_{a}=p_{b}$, and the system can be reduced to one dimension, Eqn.4.
@@
This fitness potential shows us how varying $s$ can cause a transition between a mono-stable and bistable system [park_bistability_2011](Fig.1a,b).
Because of the epistatic fitness barrier, reversing the sign of $s$ will not immediately erase the previous state, thus an evolving system will “remember” its local minimum and retain this state even if it no longer optimally fit. The red curve in Fig.1b is the characteristic hysteresis loop that is common to many history dependent systems. Fig.1e shows the degree of bistability (as quantified by the distance between wells) as a function of epistasis strength and mutation rate. Hysteresis does not occur in the white zone.
To examine the hysteresis behavior of epistatic populations, we simulated the evolution of a population of $N$ governed by a Wright-Fisher process with mutation and epistatic selection defined by Eqn.1. The cycling of the selection pressure $s$ resembles many common scenarios in nature, such as seasonal oscillations of light, temperature, nutrient and water availability [yu_distinct_2022].
From these results, we can evaluate the quality of the diffusion approximation as a function of $N$ and $u$. The diffusion approximation accurately describes the roughly linear relationship between epistatic strength and hysteresis area (Fig.1c, Eqn.S2), and holds for populations as small as $N=500$.
The setup Eqn.2 assumes that the alleles at each locus are sampled independently after selection and mutation, thus are unlinked. However, we also determine the hysteresis area when the linkage equilibrium is broken (SI.II). We find that while hysteresis area is zero for fully-linked alleles, it rapidly increases and saturates to its linkage equilibrated value, with increased recombination rate (Fig.S2).
@@
We next obtain the rate at which a population escapes over an epistatic fitness barrier. This rate measures the population’s adaptive mobility or evolvability: the likelihood of breaking free from an epistatic constraint to discover a new, fitter region of the landscape. Stronger epistasis or larger populations suppress this rate, rendering adaptation sluggish even when environmental conditions favor transition [weissman_rate_2009]. We use Eyring-Kramer theory to obtain the barrier escape rate $\kappa_{esc}\propto\sqrt{U(x)^{\prime\prime}(x_{0})|U^{\prime\prime}(x_{s})|}\exp[U(x_{0})-U(x_{s})]$, where $x_{0}$ and $x_{s}$ are the location of the minima and barrier of a generalized potential function $U(x)$ [eyring_activated_1935, kramers_brownian_1940]. Recent work has derived the correct expression for the Kramers’ rate for cases with state-dependent noise [rosas_kramers_2016], which we use to calculate the escape rate over an epistatic barrier. The scaling of the escape rate with $s,J,N,$ and $u$ is given by Eqn.5, with a negligible prefactor (SI. III).
@@
Fig.1d shows how this rate decreases exponentially with increasing population size. For large populations, barrier crossing events become exceedingly rare, in agreement with previous results on stochastic tunneling [weissman_rate_2009, altland_rare_2011], but now with explicit dependence on the selection and epistasis parameters $s$ and $J$. As the selection strength increases, escape events become much more common. These results suggest that environmental conditions and demographic constraints jointly determine whether populations remain trapped behind epistatic barriers or traverse them.
@@
To do so, we simulated the evolutionary trajectories of our epistatic two-locus two-allele population subject to a selective force with a cyclic component with amplitude $A$ and frequency $\Omega$ (to yield hysteresis) and a random Gaussian component $\eta(t)$ (that will modulate the hysteresis),
@@
We observe that noise anneals hysteresis by enabling random jumps over epistatic barriers: Epistatic hysteresis curves (2a, blue), in the presence of moderate noise (red), approach non-epistatic ones (black), indicating enhanced evolvability (Fig.2a). However, as the noise is increased further, the population is unable to keep up with an environment fluctuating too wildly, and will suffer a fitness shortfall. Fig.2b shows the sweet spot (blue markers) between evolvability and unpredictability. The fitness shortfall $\bar{\varepsilon}$ is defined as the deviation from the ideal fitness $\bar{\omega}^{*}(s(t))$ (the fitness, had there been no epistatic obstruction, i.e. with $J=0$ substituted in Eqn.3).
@@
where $\bar{\omega}(t)$ is the instantaneous mean fitness of the population at time $t$ (Eqn.3), and $T$ is the averaging time over many cycles (cf. Methods).
As expected, the optimal value of noise is a function of the ruggedness of the fitness landscape, which in our model is quantified by $J$ (Fig.2). The higher the epistatic barrier, the larger the noise the population will benefit from.
@@
Figure 3: Exploring dynamical hysteresis. A single-well potential well moving back and forth quick enough will have the effect of a double well, leading to similar hysteresis effects, remarkably, even in the absence of epistasis ($J=0$). (a) As we increase oscillation frequency from $\Omega/u=0.01$ (black) to $\Omega/u=1.0$ (red), we observe the emergence of dynamic hysteresis, and the bistability that underlies it. (b) We use the probability density of alleles (inset of (a)) to fit an effective fitness potential via Eqn.4. As we increase the cycling frequency, an effective double-well structure emerges, mimicking the role of epistasis. (c) Regardless of the true value of epistasis (black: $J_{\text{sim}}=0$, blue: $J_{\text{sim}}=0.1$, red: $J_{\text{sim}}=0.2$), we can fit an “inferred effective epistasis” $J_{\text{fit}}$ to the emergent effective potential. For $J=0$ we observe a phase transition where hysteresis emerges once the cycling rate exceeds the evolutionary relaxation rate $(\omega/u)_{\textrm{crit}}$. This transition softens and gradually vanishes with larger values of true epistasis $J_{\text{sim}}$. (d) A phase diagram for $J=0$ shows the interplay between selection noise $\sigma$ and cycling rate, and delineates the regime where dynamical hysteresis emerges.
@@
The two-locus Wright-Fisher model revealed how hysteresis emerges as a consequence of epistasis. However, it is not obvious whether hysteresis is a sensitive phenomena specific to this simple model, or it is a robust, generic property exhibited by realistic genomes where thousands of genes couple together through a complex network of epistatic interactions. For example, should we expect a large number of random (positive and negative) interactions $J_{ij}$ between gene pairs $i,j$ to “cancel out”, and wash away hysteresis? Moreover, will the physical, structural properties of a protein, which might not be simply a function of pairwise interactions, exhibit similar hysteresis loops?
We answer both questions by demonstrating the existence of hysteresis loops, first, in a complex disordered L-locus K-allele model, and then, by running biophysically realistic molecular evolution simulations, where residue–residue interactions arise from the biophysics of protein folding and structural flexibility (Fig.4).
For the L-locus K-allele model, the epistatic interaction matrix $J_{ij}^{\alpha,\beta}$ is populated from a Gaussian with mean $\langle J\rangle$ and variance $\delta J$, corresponding to increased ruggedness of the fitness landscape. The fitness potential for this more complex model is obtained through the same procedure outlined above (SI. IV),
@@
Having established that hysteresis emerges generically in both minimal and disordered epistatic systems, and out in the wild, we next ask whether the shape and size of the loop can serve as a probe to measure the evolutionary parameters of a real system. We show that by performing cyclic selection empirically, one can gain insights into $K$, $L$, $\langle J\rangle$, $\delta J$ in real systems (Fig.6).
To map the loop morphology to underlying evolutionary parameters, we expand the effective potential $\psi$ (Eqn.8) around the symmetric state $p_{li}=1/K$ in terms of the mean allelic deviation $m=\frac{1}{L}\sum_{l,i}(p_{li}-1/K)$, obtaining a coarse-grained potential $\phi(m)=a_{2}m^{2}+a_{4}m^{4}-hm$, with coefficients $a_{2}\simeq\frac{u}{K-1}\chi_{K}-\frac{1-u/K}{K-1}\lambda_{2}(J,L)+\alpha\,\delta J^{2}$ and $h\simeq\frac{1-u/K}{K-1}\lambda_{1}(L)s$, where $\chi_{K}$ arises from the entropic curvature of the mutation term and $\lambda_{1,2}$ describe the additive and pairwise contributions to $\ln\bar{\omega}(p)$. Hysteresis occurs when $a_{2}<0$, i.e., when the epistatic stabilization $\lambda_{2}(J,L)$ dominates mutation and disorder, and the barrier height between the wells approximately scales as $\Delta\phi\sim a_{2}^{2}/(4a_{4})$.
@@
In principle, the full shape of the loop, including branch curvature, asymmetry, and minor-loop closure, should contain enough information to reconstruct not only coarse descriptors such as barrier height and disorder, but also finer details of the fitness landscape and the epistatic network. Solving this inverse problem would require incorporating the complete loop trajectory rather than summary quantities such as area or squareness, as well as accounting for mutation rates, linkage, and recombination that may complicate the procedure. A promising direction could be developing analytic or machine-learning surrogates that map loop morphologies to landscape statistics. This approach could turn hysteresis measurements into a new quantitative probe of epistasis networks and fitness landscapes, but its realization lies beyond the scope of the present work.
@@ IV Methods
Here we describe the protocol used to simulate hysteresis loops for the epistatic Wright-Fisher model. Starting with a population with equal proportions of the four genotypes, we evolved the population for $1000$ generations with fixed values of $s$ and $J$. We varied $s$ in a quasi-static cycle every $1000$ generations from $0.5$ to 0.5-0.5 and back to $0.5$ to complete one evolutionary cycle. For Fig.1c, this scheme was performed independently for $N=(500,10000)$, $u=0.001$, and 20 values of $J$ linearly spaced from 0 to $0.2$. We simulated $48$ independent replicate trajectories for each parameter set $(N,J,u)$, resulting in $18000$ total trajectories for analysis. For each trajectory, we computed the area between the upper ($s\rightarrow 0^{+}$) and lower ($s\rightarrow 0^{-}$) stationary points over a full selection cycle, which we call the hysteresis area. The stationary points were defined as the average frequency of the allele under selection in the last $250$ generations of a given leg of fixed selection strength $s$. The simulations of the two-locus Wright-Fisher model for linked loci with recombination were performed by a similar quasi-static selection cycling protocol (SI. II) with $N=10000$ and $u=0.01$, for $J=0,0.1,0.2$ and a range of recombination rates r=0r=0-$0.5$. We reported the average hysteresis loop area from 48 replicate trials (Fig.S2)
@@
The simulation results presented in Fig.1d were obtain as follows. We performed simulations of a one-dimensional Ornstein-Uhlenbeck process with advection and diffusion terms given by the terms $A(p)$ and $B(p)$, respectively (Eqn.S1), which are functions of the model parameters $s,J,u$ and $N$. Starting from the local fitness maximum $p_{0}$, the stochastic variable $p(t)$ was allowed to evolve until it reached the barrier crossing point given by $p_{s}$, and the time of first crossing was recorded. We performed 1000 independent trials of this process for a range of population sizes N=500N=500-$10000$ and selection strengths $s=0.2,0.225,0.25$, with $J=0.5$. We report the average rate of escape $\kappa_{\textrm{esc}}=1/\langle T_{\textrm{esc}}\rangle$ with 95% confidence intervals determined by propagation of errors.
@@
The data presented in Fig.2a were obtained with simulation parameters $A=0.5,\Omega=0.001$, $\sigma=0,0.1371$ and $J=0,0.4$, and depict the average allele frequency of 48 independent replicate simulations. For the results presented in Fig.2b, we simulated 48 independent replicate trajectories of the epistatic Wright-Fisher model starting from equal genotype populations for model parameters J=0J=0-$0.4$, $\sigma=0$-$0.1$, u=0.0001u=0.0001-$0.01$, $N=500$, $A=0.5$, and $\Omega=0.001$. This resulted in $19200$ total trajectories for analysis. The quantity $T$ in Eqn.7 represents the total number of selection stages $s(t)$, which was chosen to complete one full hysteresis cycle for a given cycling rate $\Omega$. The results presented in Fig.3 were obtained by 48 independent replicate trajectories for the following parameters: $J=0,0.03,0.1$, $N=500$, $u=0.01$, selection amplitude $A=0.5$, cycling rate $\Omega=0.0002$-$0.02$, and selection noise $\sigma=0$-$0.1$. The resulting allele frequencies around $s=0$ were fit to the effective fitness potential model by non-linear least squares optimization using the SciPy package with Python 3.9. Uncertainty in the fit parameters were estimated as the 95% confidence interval of 500 non-parametric bootstrap simulations.
@@
We performed simulations of the $L$-locus $K$-allele Wright-Fisher model (Eqn.8) with disordered epistatis to investigate the link between epistatic interaction networks and macroscopic hysteresis in long evolutionary cycles. The simulations were performed in a similar manner as for the two-locus two-allele investigations, involving slow cycling of selection pressure with fixed epistatic interaction parameters. First, we assigned single-site fitness parameters $s_{i}^{\alpha}$ to each allele type $i$ at site $\alpha$, with $s_{i}^{\alpha}$ ranging from 0 to maximum (or minimum) fitness value $s\in(-2,2)$. The epistatic parameters $J_{ij}^{\alpha,\beta}$ were generated randomly from a Gaussian distribution $\mathcal{N}(\langle J\rangle,\delta J)$ with mean $\langle J\rangle$ and variance $\delta J$. The simulations were run in sequential stages of fixed $s$ for 1000 generations each before moving on to the next selection target. At the end of each evolutionary stage, we calculated the response of the population by Eqn.9.
@@
To simulate the evolution of protein sequences under selective pressure and epistasis, we modified the Protein Evolution Parameter Calculation (PEPC, [piszkin_extremophile_2021]) to include protein sequence flexibility prediction using the MEDUSA prediction software [vander_meersche_medusa_2021]. This method classifies amino acid residues by contribution to local structural flexibility/rigidity. Since the MEDUSA tool was trained using homologous sequences and physico-chemical properties of amino acids, it necessarily incorporates epistatic interactions between residues. Fig.4a depicts the procedure for evolving protein sequences in the modified PEPC method. Every generation, a candidate amino acid substitution is determined from a BLAST query against the Pfam database and the BLOSUM80 matrix [piszkin_extremophile_2021]. Then, a MEDUSA prediction is run on the candidate mutated sequence to assign a confidence level for the placement of each residue into one of five flexibility classes. To increase throughput, the HHblits MSA construction step of the MEDUSA prediction was performed only if the sequence had diverged by more than 80% sequence identity from its sequence at the last MSA construction, otherwise the most recent MSA was used. The predicted global flexibility $F$ of the protein sequence was calculated from the MEDUSA results by Eqn.10, where $c_{l,a}$ is the confidence that residue $l$ belongs to the MEDUSA flexibility class $a$, and $n$ is the length of the protein.
@@
Here, the classes range from least flexibility class 1 to most flexible class 5. The global flexibility of the candidate sequence is compared to the parent sequence and the mutation is accepted with a probability determined by a Metropolis-Hasting criterion defined by Eqn.11.